# Fermat near misses and the integral Hilbert property

### v1

@@ 1.1. Fermat near misses
These equations (1.1) have been recently been studied in [Eliashar2023An=4], where it was shown, rather surprisingly, that there are infinitely many solutions for $n=8$ (the case n=1n=-1 appears on Mathematics Stack Exchange and MathOverflow [Tito, MO] as a challenge problem, currently open). The solutions for $n=8$ are constructed via a complicated recursive sequence. We put this argument into a geometric context, which allows us to prove existence of infinitely many solutions for a wide list of examples. More than this, we are able to show that the integral points are not thin (see Definition 2.3); this implies for example that they are Zariski dense.
@@ 1.2. Conic bundles
A conic on a del Pezzo surface $X$ is a connected reduced curve $C\subset X$ of arithmetic genus 0 such that $K_{X}\cdot C=2$. A single conic moves in a pencil and gives rise to a conic bundle structure $\pi:X\rightarrow{\mathbb{P}}^{1}$ (see [LS22, Lem. 2.8]). Moreover if the degree of $X$ equals $1,2,4$, then there is actually a second associated conic bundle, which we call the dual of $\pi$ (see Lemma 4.1). For the surfaces in Theorem 1.1, the dual is obtained by applying the Geiser involution $w\mapstochar\rightarrow-w$. Given this, Theorem 1.1 is then a special case of the following.
@@
Assume that $X$ admits a conic bundle structure $\pi_{1}:X\rightarrow{\mathbb{P}}^{1}$ with dual $\pi_{2}:X\rightarrow{\mathbb{P}}^{1}$. Assume that $(\mathcal{X}\setminus\mathcal{D})$ admits an integral point which lies on a smooth fibre $C$ of $\pi_{1}$ which meets $D$ in a real quadratic point such that $\pi_{2}|_{D}:D\rightarrow{\mathbb{P}}^{1}$ is unramified along $C\cap D$. Then $(\mathcal{X}\setminus\mathcal{D})({\mathbb{Z}})$ is not thin.
@@
there are clearly only finitely many integral solutions. In this case the relevant boundary divisor $x^{4}+y^{4}+w^{2}=0$ has no real point. The technical unramified condition is fairly mild as it holds for all but finitely many conics (see Lemma 3.5).
@@
Assume that $X$ admits two lines and let $\pi_{1},\pi_{2}$ denote the corresponding conic bundles. Assume that $\mathcal{X}\setminus\mathcal{D}$ admits an integral point which lies on a smooth fibre $C$ of $\pi_{1}$ which meets $D$ in a real quadratic point such that $\pi_{2}|_{D}:D\rightarrow{\mathbb{P}}^{1}$ is unramified along $C\cap D$. Then $(\mathcal{X}\setminus\mathcal{D})({\mathbb{Z}})$ is Zariski dense.
@@
Let $k$ be a number field and $X$ be a smooth projective surface over $k$ which admits two distinct conic bundle structures $\pi_{i}:X\rightarrow{\mathbb{P}}^{1}$ for $i=1,2$. Let $D\subset X$ be a smooth irreducible divisor. Let $S$ be a finite set of places containing the archimedean places and $\mathcal{X}$ a projective model for $X$ over ${\mathcal{O}}_{k,S}$ and $\mathcal{D}$ the closure of $D$ in $\mathcal{X}$. Assume that the following conditions hold.
@@
      The map $\pi_{2}|_{D}:D\rightarrow{\mathbb{P}}^{1}$ is unramified along $C\cap D$.
@@
An interesting challenge case is when n-n is a square; here there are obvious solutions, but infinitude is not known in any case aside from when $n=-4m^{4}$. Outside this case, there is no conic bundle structure by Proposition 5.1.
@@ 1.3. Method of proof and related literature
Theorem 1.5 is proved via the double fibration method. This method was first introduced by Corvaja and Zannier [Corvaja2017OnVarieties] to show that rational points on the Fermat quartic surface are not thin. This method was subsequently generalised by Demeio [Demeio2020Non-rationalProperty, Demeio2021EllipticProperty] (elliptic surfaces) and Streeter [Streeter2021HilbertVarieties] (conic bundles).
The closest work to ours is Coccia [Coccia2019TheSurfaces, Coccia2024ASurfaces]. Coccia proves in [Coccia2024ASurfaces, Thm. 1.7] a result on the integral Hilbert property for double conic bundle surfaces, with possibly singular boundary divisor. He uses this to obtain the integral Hilbert property [Coccia2024ASurfaces, Thm. 1.6] for del Pezzo surfaces after a possible finite field extension and after enlarging the set of places. Some differences with our Theorem 1.5 are that Coccia assumes in [Coccia2024ASurfaces, Thm. 1.7 (H2)] that there is a non-thin set of fibres which have a Zariski dense set of integral points, whereas our main applications only require finding one “good” integral point. Moreover the assumptions in [Coccia2024ASurfaces, Thm. 1.7] imply that the fibres of the two conic fibrations have intersection number $2$, so this does not cover the case of del Pezzo surfaces of degree $2$ as in (1.1), where the intersection number is $4$.
Our main innovation is to make this method work with weaker assumptions, and with particular care constructing affine conics over $\mathbb{Z}$} with infinitely many integral points, since this is more difficult over $\mathbb{Q}$} as there is only one place at infinity (see Lemma 2.5). We also require a result on strong approximation for integral points on conics (Proposition 2.7), which we combine with a result on lifting integral points to covers (Lemma 3.4) and topological arguments in Lemma 3.6 to generate conics in the family with infinitely many integral points. We keep running our process to show that there are non-thin many conics with infinitely many integral points, from where our proof follows a similar strategy to that of Coccia.
In private communication, Coccia has informed us of forthcoming work where he applies the double fibration method to prove the integral Hilbert property for some affine cubic surfaces over $\mathbb{Z}$} with possibly singular boundary divisor, including important classical surfaces such as Markoff type cubic surfaces. His approach builds upon a sketch of a proof, given in [Coccia2024ASurfaces, §6.4], regarding how to apply his method to a specific affine cubic surface over $\mathbb{Z}$}.
@@ Acknowledgements
Both authors were supported by UKRI Future Leaders Fellowship MR/V021362/1. We are grateful to Julian Demeio and Harry Shaw for useful discussions, Sam Streeter for detailed comments which led to Lemma 4.4. We thank Kumayl Vanat for providing the data in Table 1.1, and Ari Shnidman for making us aware of the paper [Eliashar2023An=4] and discussing the geometry of the surface. We also thank Simone Coccia for various comments and pointing out a mistake in an earlier version of Lemma 3.5. JA thanks the Max Planck Institute for Mathematics in Bonn for their hospitality and financial support. JA is member of the INdAM group GNSAGA.
@@ 2.1. Integral Hilbert property
Let $X$ be an integral variety over a field $k$. A thin map is a morphism $Y\rightarrow X$ from an integral variety which admits no rational section. A thin subset of $X(k)$ is a subset contained in the images of the rational points of the union of finitely many thin maps.
@@
An $S$-integral point of $\mathcal{X}\setminus\mathcal{D}$ is a section $s_{P}$ of $\mathcal{X}\rightarrow\mathrm{Spec}({\mathcal{O}}_{k,S})$ avoiding $\mathcal{D}$, i.e. for each $\mathfrak{p}\in\mathrm{Spec}({\mathcal{O}}_{k,S})$ we have that $s_{P}(\mathfrak{p})\notin\mathcal{D}_{\mathfrak{p}}$. We denote the set of these points by $(\mathcal{X}\setminus\mathcal{D})({\mathcal{O}}_{k,S})$. If $D$ and the model is clear from the context, we also denote by $X^{\circ}=X\setminus D$ and $X^{\circ}({\mathcal{O}}_{k,S})=(\mathcal{X}\setminus\mathcal{D})({\mathcal{O}}_{k,S})$.
@@ 2.2. Integral points on curves
We now take the following set up: let $C$ be a smooth projective curve of genus 0 over $k$ and $D$ a reduced divisor on $C$ of degree at most $2$; we often informally refer to $D$ as the “points at infinity”. Let $S$ be a finite set of places of $k$ containing the archimedean ones and $\mathcal{C}$ a normal projective model for $C$ over ${\mathcal{O}}_{k,S}$, with $\mathcal{D}$ the closure of $D$ in $\mathcal{C}$. We write $C^{\circ}=C\setminus D$ and $C^{\circ}({\mathcal{O}}_{k,S})=(\mathcal{C}\setminus\mathcal{D})({\mathcal{O}}_{k,S})$.
@@
Another result we require shows that only finitely many integral points lift to certain covers [Coccia2024ASurfaces, Lem. 5.3], as a consequence of Siegel’s Theorem.
@@
Assume that $\deg D=2$. Let $V$ be a smooth projective irreducible curve and $f:V\rightarrow C$ be a finite map. Assume that one of the following holds:
@@
$\deg_{v}:T(k_{v})\rightarrow\operatorname{Hom}(\widehat{T}^{\Gamma_{v}},{\mathbb{Z}}),\quad t_{v}\mapstochar\rightarrow\left(\,\chi_{v}\mapstochar\rightarrow v(\chi_{v}(t_{v}))\,\right)$
@@
$0\rightarrow T({\mathcal{O}}_{v})\rightarrow T(k_{v})\rightarrow\operatorname{Hom}(\widehat{T}^{\Gamma_{v}},{\mathbb{Z}})\rightarrow 0.$ (2.1)
@@ 2.3. Del Pezzo surfaces
where $\deg F=4$, provided the characteristic of $k$ is not equal to $2$. The anticanonical map $X\rightarrow{\mathbb{P}}^{2}$ is given by projecting to $(x,y,z)$, and realises the surface as a double cover of ${\mathbb{P}}^{2}$ ramified in the smooth quartic curve $F(x,y,z)=0$. This induces the Geiser involution $\iota:X\rightarrow X$ which swaps the two sheets of the cover.
@@ 2.4. Conic bundles
Let $X$ be a smooth projective surface $X$ over a field $k$. A conic bundle is a dominant morphism $\pi:X\rightarrow\mathbb{P}^{1}$ such that the fibres are isomorphic to plane conics.
@@
Let $\pi:X\rightarrow{\mathbb{P}}^{1}$ be a conic bundle surface with fibre class $C$. Then
@@
This is a special case of [Frei2018RationalSurfaces, Prop. 5.2]. We sketch a proof, since it also furnishes an effective method to construct a conic bundle structure explicitly. Moreover the Stein factorisation step at the end of loc. cit. is actually not required.
@@
One achieves an explicit Galois descent and construction of the map as follows. We have $\dim\mathrm{H}^{0}(X,\mathfrak{L})=2^{n}$. The image of the induced map $X\rightarrow{\mathbb{P}}^{2^{n}-1}$ is a curve of genus 0. As $X(k)\neq\emptyset$ this curve has a rational point. Thus choosing an isomorphism of this curve with ${\mathbb{P}}^{1}$ yields a map $X\rightarrow{\mathbb{P}}^{1}$ which is the required conic bundle. ∎
@@ 3.1. Fundamental groups
Let $X$ be a smooth proper surface over $k$ and $D\subset X$ an effective divisor. Let $E\subset X$ be a collection of pairwise disjoint (1)(-1)-curves and $f:X\rightarrow X^{\prime}$ the blow down of $E$. Then $\pi_{1}(X\setminus(D\cup E))\cong\pi_{1}(X^{\prime}\setminus f(D))$.
@@
Let $X$ be a smooth proper surface over $k$ which admits two conic bundle structures $\pi_{i}:X\rightarrow{\mathbb{P}}^{1}$ with fibre classes $C_{1},C_{2}$, respectively, such that $C_{1}\cdot C_{2}=1$. Let $E$ be the divisor of curves which are constant with respect to both fibrations and $D\subset X$ an integral divisor with $C_{1}\cdot D=C_{2}\cdot D=2$. Then $X\setminus(D\cup E)$ is not simply connected.
@@
We may assume that $k$ is algebraically closed. Blowing down fibre components of $\pi_{1}$, we obtain a map $f:X\rightarrow X^{\prime}$ to a smooth projective surface which admits a conic bundle with smooth fibres, i.e. a ruling. We now carefully analyse the possibilities for $X^{\prime}$.
@@ 3.2. Preparations
      The map $\pi_{2}|_{D}:D\rightarrow{\mathbb{P}}^{1}$ is unramified along $C\cap D$.
@@
Let $\Omega\subset{\mathbb{P}}^{1}$ be a finite collection of closed points. Then for all but finitely many $P\in{\mathbb{P}}^{1}(k)$, the branch locus of the map $\pi_{2}:C_{1,P}\rightarrow{\mathbb{P}}^{1}$, induced by $\pi_{2}$, is disjoint from $\Omega$.
@@
If $\pi_{2}:C_{1,P}\rightarrow{\mathbb{P}}^{1}$ ramifies over a point $Q\in{\mathbb{P}}^{1}$, then $C_{1,P}\cap C_{2,Q}$ is non-reduced. However, by Lemma 2.10, the conic $C_{2,Q}\subset X$ is reduced. Therefore, as $k$ has characteristic 0, the restriction $\pi_{1}:C_{2,Q}\rightarrow{\mathbb{P}}^{1}$ is generically smooth, which implies that $C_{1,P}\cap C_{2,Q}$ is non-reduced for only finitely many $P$. The result now easily follows. ∎
@@
Let $f:Y\rightarrow{\mathbb{P}}^{1}$ be a finite morphism with $Y$ a smooth projective curve and $\deg f>1$. Let $P\in{\mathbb{P}}^{1}(k)$ and $C:=\pi_{1}^{-1}(P)$. The set $\pi_{2}(C^{\circ}({\mathcal{O}}_{k,S}))\cap f(Y(k))\subset{\mathbb{P}}^{1}(k)$ is finite if
@@
    the scheme $C\cap D$ is reduced and the maps $f:Y\rightarrow{\mathbb{P}}^{1}$ and $\pi_{2}|_{C}:C\rightarrow{\mathbb{P}}^{1}$ have disjoint branch loci.
@@
Let $\mathcal{Y}\rightarrow{\mathbb{P}}^{1}$ be a model for $Y$. Consider the fibre product
@@
Let $\Omega\subset D$ be a finite collection of closed points. Then for all but finitely many $P\in{\mathbb{P}}^{1}(k)$ we have $C_{1,P}\cap\Omega=\emptyset$. In particular, for all but finitely many $P\in{\mathbb{P}}^{1}(k)$ the map $\pi_{2}|_{D}:D\rightarrow{\mathbb{P}}^{1}$ is unramified along $C_{1,P}\cap D$.
@@
Immediate from the fact that the linear system given by the conics is base-point free and that $\pi_{2}|_{D}:D\rightarrow{\mathbb{P}}^{1}$ ramifies in only finitely many points. ∎
@@
Our version of Zariski density is as follows.
@@
Let $C$ be a smooth conic as in hypothesis (3) of Theorem 1.5. By Lemma 2.5, there is a place $v\in S$ such that $(C\cap D)(k_{v})\neq\emptyset$. Let $W_{D}\subseteq D(k_{v})$ be the complement of the ramification points of $\pi_{2}|_{D}:D\rightarrow{\mathbb{P}}^{1}$; this is open in $D(k_{v})$ as $D$ is smooth. Hence $\pi_{2}(W_{D})\subset{\mathbb{P}}^{1}(k_{v})$ is open as a smooth morphism is open for the $v$-adic topology. We now consider the cylinder $W:=\pi_{2}^{-1}(\pi_{2}(W_{D}))\subset X(k_{v})$, which is an open subset of $X(k_{v})$. For any $P\in\pi_{2}(W)$, since $\pi_{2}(W)=\pi_{2}(W_{D})$ we have $(C_{2,P}\cap D)(k_{v})\neq\emptyset$.
@@
As $T$ is thin, there is a finite collection $\phi_{i}:Y_{i}\rightarrow\mathbb{P}^{1}$, $i=1,\ldots,r$, with $Y_{i}$ a smooth geometrically irreducible curve and $\deg(\phi_{i})>1$, such that $T\subseteq\bigcup_{i=1}^{r}\phi_{i}(Y_{i}(k))$. The result then follows from applying Lemma 3.4 to each $\phi_{i}$. ∎
@@
We say that a cover $\phi:Y\rightarrow X$ is $\pi$–unramified if:
@@
We argue by contradiction. Assume that there exist finitely many finite maps $\phi_{i}:Y_{i}\rightarrow X$ of degree >1>1, with $Y_{i}$ a geometrically irreducible normal projective variety, such that $X^{\circ}({\mathcal{O}}_{k,S})\setminus\cup_{i}(\phi_{i}(Y_{i}(k))$ is contained in a curve $Z$. By Zariski’s purity theorem the branch locus $B_{i}$ of $\phi_{i}$ is a divisor.
@@
Not all conic bundles with Zariski dense set of integral points satisfy the conclusion of Lemma 3.6. Consider the affine surface $U:xy=t$ over $\mathbb{Z}$}. This is isomorphic to $\mathbb{A}^{2}_{{\mathbb{Z}}}$, hence $U({\mathbb{Z}})$ is not thin. It is equipped with various conic bundles, e.g. the fibration $f:U\rightarrow\mathbb{A}^{1},(x,y,t)\rightarrow t$. But each fibre of $f$ has only finitely many integral points. This shows that our assumptions are necessary for our method (here the boundary divisor is reducible).

### v2

@@ 1.1. Fermat near misses
These equations (1.1) have been recently been studied in [Eliashar2023An=4], where it was shown, rather surprisingly, that there are infinitely many solutions for $n=8$ (the case n=1n=-1 appears on Mathematics Stack Exchange and MathOverflow [Tito, MO] as a challenge problem, currently open). The solutions for $n=8$ are constructed in [Eliashar2023An=4] via a complicated recursive sequence. We put this argument into a geometric context, which allows us to prove existence of infinitely many solutions for a wide list of examples. More than this, we are able to show that the integral points are not thin (see Definition 2.3); this implies for example that they are Zariski dense.
@@ 1.2. Conic bundles
A conic on a del Pezzo surface $X$ is a connected reduced curve $C\subset X$ of arithmetic genus 0 such that $K_{X}\cdot C=2$. A single conic moves in a pencil and gives rise to a conic bundle structure $\pi:X\to{\mathbb{P}}^{1}$ (see [LS22, Lem. 2.8]). Moreover if the degree of $X$ equals $1,2,4$, then there is actually a second associated conic bundle, which we call the dual of $\pi$ (see Lemma 4.1). For the surfaces in Theorem 1.1, the dual is obtained by applying the Geiser involution $w\mapsto-w$. Given this, Theorem 1.1 is then a special case of the following.
@@
Assume that $X$ admits a conic bundle structure $\pi_{1}:X\to{\mathbb{P}}^{1}$ with dual $\pi_{2}:X\to{\mathbb{P}}^{1}$. Assume that $(\mathcal{X}\setminus\mathcal{D})$ admits an integral point which lies on a smooth fibre $C$ of $\pi_{1}$ which meets $D$ in a real quadratic point such that $\pi_{2}|_{D}:D\to{\mathbb{P}}^{1}$ is unramified along $C\cap D$. Then $(\mathcal{X}\setminus\mathcal{D})({\mathbb{Z}})$ is not thin.
@@
there are clearly only finitely many integral solutions. In this case the relevant boundary divisor $x^{4}+y^{4}+w^{2}=0$ has no real point. The technical unramified condition is fairly mild as it holds for all but finitely many conics (see Lemma 3.5); in fact it is so rare that it is automatically satisfied in the situation of Theorem 1.1 (see Lemma 5.4).
@@
Assume that $X$ admits two lines and let $\pi_{1},\pi_{2}$ denote the corresponding conic bundles. Assume that $\mathcal{X}\setminus\mathcal{D}$ admits an integral point which lies on a smooth fibre $C$ of $\pi_{1}$ which meets $D$ in a real quadratic point such that $\pi_{2}|_{D}:D\to{\mathbb{P}}^{1}$ is unramified along $C\cap D$. Then $(\mathcal{X}\setminus\mathcal{D})({\mathbb{Z}})$ is Zariski dense.
@@
Let $k$ be a number field and $X$ be a smooth projective surface over $k$ which admits two distinct conic bundle structures $\pi_{i}:X\to{\mathbb{P}}^{1}$ for $i=1,2$. Let $D\subset X$ be a smooth irreducible divisor. Let $S$ be a finite set of places containing the archimedean places and $\mathcal{X}$ a projective model for $X$ over ${\mathcal{O}}_{k,S}$ and $\mathcal{D}$ the closure of $D$ in $\mathcal{X}$. Assume that the following conditions hold.
@@
      The map $\pi_{2}|_{D}:D\to{\mathbb{P}}^{1}$ is unramified along $C\cap D$.
@@
An interesting case is when n-n is a square; here there are obvious solutions, but infinitude is not known in any case aside from when $n=-4m^{4}$ by Theorem 1.1. Outside this case, there is no conic bundle structure by Proposition 5.1.
@@ 1.3. Method of proof and related literature
Theorem 1.5 is proved via the double fibration method. This method was first introduced by Corvaja and Zannier [Corvaja2017OnVarieties] to show that rational points on the Fermat quartic surface are not thin. This method was subsequently generalised by Demeio [Demeio2020Non-rationalProperty, Demeio2021EllipticProperty] (elliptic surfaces) and Streeter [Streeter2021HilbertVarieties] (conic bundles). Conic bundles and infinite automorphism groups are also used in [KolVP, Kolllar] to prove Zariski density of integral points on certain affine cubic surfaces with singular boundary divisor.
The closest work to ours is Coccia [Coccia2019TheSurfaces, Coccia2024ASurfaces]. Coccia proves in [Coccia2024ASurfaces, Thm. 1.7] a result on the integral Hilbert property for double conic bundle surfaces, with possibly singular boundary divisor. He uses this to obtain the integral Hilbert property [Coccia2024ASurfaces, Thm. 1.6] for del Pezzo surfaces after a possible finite field extension and after enlarging the set of places. A sketch of a proof is also given in [Coccia2024ASurfaces, §6.4] regarding how to apply his method to a specific affine cubic surface over $\mathbb{Z}$}. The assumptions in [Coccia2024ASurfaces, Thm. 1.7] imply that the fibres of the two conic fibrations have intersection number $2$, so this does not cover the case of del Pezzo surfaces of degree $2$ as in (1.1), where the intersection number is $4$.
Our main innovation is to make this method work with weaker assumptions, and with particular care constructing affine conics over $\mathbb{Z}$} with infinitely many integral points, since this is more difficult over $\mathbb{Q}$} as there is only one place at infinity (see Lemma 2.5). We also require a result on strong approximation for integral points on conics (Proposition 2.7), which we combine with a result on lifting integral points to covers (Lemma 3.4) and topological arguments in Lemma 3.6 to generate conics in the family with infinitely many integral points. We keep running our process to show that there are non-thin many conics with infinitely many integral points, from where our proof follows a similar strategy to [Coccia2024ASurfaces, Thm. 1.7].
In private communication, Coccia has informed us of forthcoming work where he applies the double fibration method to prove the integral Hilbert property for some affine cubic surfaces over $\mathbb{Z}$} with possibly singular boundary divisor, including important classical surfaces such as Markoff type cubic surfaces. His approach uses the method sketched in [Coccia2024ASurfaces, §6.4], which is a similar strategy to our own.
@@
Funding
This work was supported by UKRI Future Leaders Fellowship MR/V021362/1. JA thanks the Max Planck Institute for Mathematics in Bonn for their hospitality and financial support. JA is member of Italian “National Group for Algebraic and Geometric Structures, and their Application” (GNSAGA - INdAM).
@@ Acknowledgements
We are grateful to Julian Demeio and Harry Shaw for useful discussions, Sam Streeter for detailed comments which led to Lemma 4.4. We thank Kumayl Vanat for providing the data in Table 1.1, and Ari Shnidman for making us aware of the paper [Eliashar2023An=4] and discussing the geometry of the surface. We also thank Simone Coccia for various comments and pointing out a mistake in an earlier version of Lemma 3.5.
@@ 2.1. Integral Hilbert property
We recall some basic facts about thin sets and the integral Hilbert property here. For additional background and context, see the recent survery article [FehmJav2025].
@@
Let $X$ be an integral variety over a field $k$. A thin map is a morphism $Y\to X$ from an integral variety which admits no rational section. A thin subset of $X(k)$ is a subset contained in the images of the rational points of the union of finitely many thin maps.
@@
An $S$-integral point of $\mathcal{X}\setminus\mathcal{D}$ is a section $s_{P}$ of $\mathcal{X}\to\mathrm{Spec}({\mathcal{O}}_{k,S})$ avoiding $\mathcal{D}$, i.e. for each $\mathfrak{p}\in\mathrm{Spec}({\mathcal{O}}_{k,S})$ we have that $s_{P}(\mathfrak{p})\notin\mathcal{D}_{\mathfrak{p}}$. We denote the set of these points by $(\mathcal{X}\setminus\mathcal{D})({\mathcal{O}}_{k,S})$. If $D$ and the model is clear from the context, we also denote by $X^{\circ}=X\setminus D$ and $X^{\circ}({\mathcal{O}}_{k,S})=(\mathcal{X}\setminus\mathcal{D})({\mathcal{O}}_{k,S})$.
@@ 2.2. Integral points on curves
We now take the following set up: let $C$ be a smooth projective curve of genus 0 over $k$ and $D$ a reduced effective divisor on $C$ of degree at most $2$; we informally refer to $D$ as the “points at infinity”. Let $S$ be a finite set of places of $k$ containing the archimedean ones and $\mathcal{C}$ a normal projective model for $C$ over ${\mathcal{O}}_{k,S}$, with $\mathcal{D}$ the closure of $D$ in $\mathcal{C}$. We write $C^{\circ}=C\setminus D$ and $C^{\circ}({\mathcal{O}}_{k,S})=(\mathcal{C}\setminus\mathcal{D})({\mathcal{O}}_{k,S})$.
@@
Another result [Coccia2024ASurfaces, Lem. 5.3] we require shows that only finitely many integral points lift to certain covers, as a consequence of Siegel’s Theorem.
@@
Assume that $\deg D=2$. Let $V$ be a smooth projective irreducible curve and $f:V\to C$ be a finite map. Assume that one of the following holds:
@@
$\deg_{v}:T(k_{v})\to\operatorname{Hom}(\widehat{T}^{\Gamma_{v}},{\mathbb{Z}}),\quad t_{v}\mapsto\left(\,\chi_{v}\mapsto v(\chi_{v}(t_{v}))\,\right)$
@@
$0\to T({\mathcal{O}}_{v})\to T(k_{v})\to\operatorname{Hom}(\widehat{T}^{\Gamma_{v}},{\mathbb{Z}})\to 0.$ (2.1)
@@ 2.3. Del Pezzo surfaces
where $\deg F=4$, provided the characteristic of $k$ is not equal to $2$. The anticanonical map $X\to{\mathbb{P}}^{2}$ is given by projecting to $(x,y,z)$, and realises the surface as a double cover of ${\mathbb{P}}^{2}$ ramified in the smooth quartic curve $F(x,y,z)=0$. This induces the Geiser involution $\iota:X\to X$ which swaps the two sheets of the cover.
@@ 2.4. Conic bundles
Let $X$ be a smooth projective surface $X$ over a field $k$. A conic bundle is a morphism $\pi:X\to\mathbb{P}^{1}$ whose fibres are isomorphic to plane conics.
@@
Let $\pi:X\to{\mathbb{P}}^{1}$ be a conic bundle surface with fibre class $C$. Then
@@
This is a special case of [Frei2018RationalSurfaces, Prop. 5.2]. We sketch a proof, since it also furnishes an effective method to construct a conic bundle structure explicitly. Moreover the Stein factorisation step at the end of loc. cit. is not actually required.
@@
One achieves an explicit Galois descent and construction of the map as follows. We have $\dim\mathrm{H}^{0}(X,\mathfrak{L})=2^{n}$. The image of the induced map $X\to{\mathbb{P}}^{2^{n}-1}$ is a curve of genus 0. As $X(k)\neq\emptyset$ this curve has a rational point. Thus choosing an isomorphism of this curve with ${\mathbb{P}}^{1}$ yields a map $X\to{\mathbb{P}}^{1}$ which is the required conic bundle. ∎
@@ 3.1. Fundamental groups
Let $X$ be a smooth proper surface over $k$ and $D\subset X$ an effective divisor. Let $E\subset X$ be a collection of pairwise disjoint (1)(-1)-curves and $f:X\to X^{\prime}$ the blow down of $E$. Then $\pi_{1}(X\setminus(D\cup E))\cong\pi_{1}(X^{\prime}\setminus f(D))$.
@@
Let $X$ be a smooth proper surface over $k$ which admits two conic bundle structures $\pi_{i}:X\to{\mathbb{P}}^{1}$ with fibre classes $C_{1},C_{2}$, respectively, such that $C_{1}\cdot C_{2}=1$. Let $E$ be the divisor of curves which are constant with respect to both fibrations and $D\subset X$ an integral divisor with $C_{1}\cdot D=C_{2}\cdot D=2$. Then $X\setminus(D\cup E)$ is not simply connected.
@@
We may assume that $k$ is algebraically closed. Blowing down fibre components of $\pi_{1}$, we obtain a map $f:X\to X^{\prime}$ to a smooth projective surface which admits a conic bundle with smooth fibres, i.e. a ruling. We now carefully analyse the possibilities for $X^{\prime}$.
@@ 3.2. Preparations
      The map $\pi_{2}|_{D}:D\to{\mathbb{P}}^{1}$ is unramified along $C\cap D$.
@@
Let $\Omega\subset{\mathbb{P}}^{1}$ be a finite collection of closed points. Then for all but finitely many $P\in{\mathbb{P}}^{1}(k)$, the branch locus of the map $\pi_{2}:C_{1,P}\to{\mathbb{P}}^{1}$, induced by $\pi_{2}$, is disjoint from $\Omega$.
@@
If $\pi_{2}:C_{1,P}\to{\mathbb{P}}^{1}$ ramifies over a point $Q\in{\mathbb{P}}^{1}$, then $C_{1,P}\cap C_{2,Q}$ is non-reduced. However, by Lemma 2.10, the conic $C_{2,Q}\subset X$ is reduced. Therefore, as $k$ has characteristic 0, the restriction $\pi_{1}:C_{2,Q}\to{\mathbb{P}}^{1}$ is generically smooth, which implies that $C_{1,P}\cap C_{2,Q}$ is non-reduced for only finitely many $P$. The result now easily follows. ∎
@@
Let $f:Y\to{\mathbb{P}}^{1}$ be a finite morphism with $Y$ a smooth projective curve and $\deg f>1$. Let $P\in{\mathbb{P}}^{1}(k)$ and $C:=\pi_{1}^{-1}(P)$. The set $\pi_{2}(C^{\circ}({\mathcal{O}}_{k,S}))\cap f(Y(k))\subset{\mathbb{P}}^{1}(k)$ is finite if
@@
    the scheme $C\cap D$ is reduced and the maps $f:Y\to{\mathbb{P}}^{1}$ and $\pi_{2}|_{C}:C\to{\mathbb{P}}^{1}$ have disjoint branch loci.
@@
Let $\mathcal{Y}\to{\mathbb{P}}^{1}$ be a model for $Y$. Consider the fibre product
@@
Let $\Omega\subset D$ be a finite collection of closed points. Then for all but finitely many $P\in{\mathbb{P}}^{1}(k)$ we have $C_{1,P}\cap\Omega=\emptyset$. In particular, for all but finitely many $P\in{\mathbb{P}}^{1}(k)$ the map $\pi_{2}|_{D}:D\to{\mathbb{P}}^{1}$ is unramified along $C_{1,P}\cap D$.
@@
Immediate from the fact that the linear system given by the conics is base-point free and that $\pi_{2}|_{D}:D\to{\mathbb{P}}^{1}$ ramifies in only finitely many points. ∎
@@
Our strong version of Zariski density is as follows.
@@
Let $C$ be a smooth conic as in hypothesis (3) of Theorem 1.5. By Lemma 2.5, there is a place $v\in S$ such that $(C\cap D)(k_{v})\neq\emptyset$. Let $W_{D}\subseteq D(k_{v})$ be the complement of the ramification points of $\pi_{2}|_{D}:D\to{\mathbb{P}}^{1}$; this is open in $D(k_{v})$ as $D$ is smooth. Hence $\pi_{2}(W_{D})\subset{\mathbb{P}}^{1}(k_{v})$ is open as an étale morphism is open for the $v$-adic topology. We now consider the cylinder $W:=\pi_{2}^{-1}(\pi_{2}(W_{D}))\subset X(k_{v})$, which is an open subset of $X(k_{v})$. For any $P\in\pi_{2}(W)$, since $\pi_{2}(W)=\pi_{2}(W_{D})$ we have $(C_{2,P}\cap D)(k_{v})\neq\emptyset$.
@@
As $T$ is thin, there is a finite collection $\phi_{i}:Y_{i}\to\mathbb{P}^{1}$, $i=1,\ldots,r$, with $Y_{i}$ a smooth geometrically irreducible curve and $\deg(\phi_{i})>1$, such that $T\subseteq\bigcup_{i=1}^{r}\phi_{i}(Y_{i}(k))$. The result then follows from applying Lemma 3.4 to each $\phi_{i}$. ∎
@@
We say that a cover $\phi:Y\to X$ is $\pi$–unramified if:
@@
We argue by contradiction. Assume that there exist finitely many finite maps $\phi_{i}:Y_{i}\to X$ of degree >1>1, with $Y_{i}$ a geometrically irreducible normal projective variety, such that $X^{\circ}({\mathcal{O}}_{k,S})\setminus\cup_{i}(\phi_{i}(Y_{i}(k))$ is contained in a curve $Z$. By Zariski’s purity theorem the branch locus $B_{i}$ of $\phi_{i}$ is a divisor.
@@
Not all conic bundles with Zariski dense set of integral points satisfy the conclusion of Lemma 3.6. Consider the affine surface $U:xy=t$ over $\mathbb{Z}$}. This is isomorphic to $\mathbb{A}^{2}_{{\mathbb{Z}}}$, hence $U({\mathbb{Z}})$ is not thin. It is equipped with various conic bundles, e.g. the fibration $f:U\to\mathbb{A}^{1},(x,y,t)\to t$. But each fibre of $f$ has only finitely many integral points. This shows that our assumptions are necessary for our method (here the boundary divisor is reducible).